The cube is immersed in one liquid by 15 cm and in the other by 40 cm. To what depth will it sink into a liquid whose density is the arithmetic mean of the densities of the first two liquids?
1
Expert's answer
2020-06-09T13:19:33-0400

Newton's second law for all three cases:

"\begin{cases}
mg=\rho_1gh_1\\
mg=\rho_2gh_2\\
mg=\frac{\rho_1+\rho_2}{2}gH
\end{cases}"

we express H and density from these equations:

"\begin{cases}
\rho_1 = \frac{mg}{gh_1}\\
\rho_2 = \frac{mg}{gh_2}\\
H=\frac{2mg}{g(\rho_1+\rho_2)}
\end{cases}"

then

"H= \frac{2mg}{g(\frac{mg}{gh_1}+\frac{mg}{gh_2})}=\frac{2h_1h_2}{h_1+h_2} = \frac{2\cdot15\cdot40}{15+40} = 54.55 cm"